Tohban Report 2013-03-06

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 27 Feb 2013
End Date: 6 Mar 2013
Tohban: Lindsay Glesener
Tohban email:
Next Tohban: Juan Carlos Martinez Oliveros
List all reports


Solar Activity

The Sun remained fairly quiet this week, with only a few C flares and 1 M class flare. There are 10 AR on the disk right now, pretty well spread across longitude. The region that produced the C and M flares (11686) will stick around for a couple more days before rotating off the disk. STEREO-B doesn't show anything too exciting about to rotate on.

[UPDATE] Following up on Mark's comment, it looks like there was a largish backside CME on March 5th. My guess at its position indicates that it was not from a region close to rotating on-disk. Even so, perhaps it will give us some more large events which could be observed partly-occulted from Earth's POV.

How many GOES flares occurred?

 Flares above B, C, M, X class were     18    3     1     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were     12     2     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      6     0     0     0

There were RHESSI flares/GOES flares 153 / 22 over the time range 27-Feb-13 06-Mar-13

Memory Management

The SSR is emptying after each pass set.

Spacecraft Status

CP2 is at ~110K, not a big change from last week.

On Day 64 the EDAC (error detection and correction circuitry) counter started increasing, indicating an error on the onboard CPU. This is not so unusual; it happens from time to time. The error was corrected and the error counter cleared. The error counter was increasing for about a half hour. No data was lost.

Data Gaps

No data gaps noticed by the tohban. According to Jeremy, there are some very small gaps (<4 sec) on DOY 64 and 65. Those will be replayed today.

Detector issues

There was an isolated spike in Detector 8 on March 2 at ~15:48 UT. This was during a flare, just after exiting the SAA, so maybe a particularly stressful time for that detector. No other spikes or problems were noticed for the week.

We didn't do anything with Detector 2 this week. The plan is to stay at this HV and threshold setting for at least another week while we assess its performance.

Other notes

Decimation Normal/vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/- 4 minutes
Night time data (rears) Taking data at nights
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? No problems.
Personal tools